I need to know how to put this word problem into an algebraic expression. A supply store sold a wirebound notebook for $2.50 and a college-ruled notebook for $2.30. There were a combination of 50 notebooks sold for a total of $118.60. How many of each type were sold?
Answer by rfer(16322)

w+c=50 c=50-w
2.5w+2.3c=118.6
2.5w+2.3(50-w)=118.6
2.5w+115-2.3w=118.6
.2w=3.6
w=18
--------------
c=50-18
c=32
-------------
32*2.3+18*2.5=
73.6+45=$118.60